# MoLA: Motion Generation and Editing with Latent Diffusion Enhanced by Adversarial Training

### v2

@@ Abstract.
In motion generation, controllability as well as generation quality and speed is becoming more and more important. There are various motion editing tasks, such as in-betweening, upper body editing, and path-following, but existing methods perform motion editing with a data-space diffusion model, which is slow in inference compared to a latent diffusion model. In this paper, we propose MoLA, which provides fast and high-quality motion generation and also can deal with multiple editing tasks in a single framework. For high-quality and fast generation, we employ a variational autoencoder and latent diffusion model, and improve the performance with adversarial training. In addition, we apply a training-free guided generation framework to achieve various editing tasks with motion control inputs. We quantitatively show the effectiveness of adversarial learning in text-to-motion generation, and demonstrate the applicability of our editing framework to multiple editing tasks in the motion domain. We will make our code publicly available.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Computer visionccs: Computing methodologies Motion processing
Figure 1. MoLA achieves fast and high-quality human motion generation given textual descriptions while enabling motion editing applications. With MoLA, we can deal with various types of motion editing tasks in a single framework. The above images show the standard text-to-motion generation (as a no-control input case) and three motion editing tasks (with control inputs): the control input for the start-end positions and motion in-between task (left), for the lower body positions and upper body editing task (center), for the motion trajectory and path-following generation task (right).
@@ 1. Introduction
Fast generation
High-quality generation
Training-free editing
MDM-based methods (Tevet et al., 2023; Karunratanakul et al., 2023b; Xie et al., 2023; Karunratanakul et al., 2023a) - -
Motion Latent Diffusion (MLD) (Chen et al., 2023) - -
VQ-based methods (Zhang et al., 2023b; Guo et al., 2023; Lou et al., 2023; Pinyoanuntapong et al., 2024)   -
MoLA (ours)
Table 1. Comparison of three properties satisfied by MoLA and other text-to-motion methods. $\dagger$ Some editing tasks can be done in a training-free manner, but there are editing tasks that cannot be done.
Motion generation is a fundamental task in multimedia production and computer animation, and it can be challenging due to the diversity of possible motions and the difficulty and cost of obtaining high-quality data. Generating human motions from textual descriptions, i.e., text-to-motion generation, is an emerging task that may play an important role in a broad range of applications such as games, film, the metaverse, and virtual reality. Over the past few years, text-to-motion has generated intensive research interest (Tevet et al., 2023; Wang et al., 2023; Zhong et al., 2023; Guo et al., 2023; Lou et al., 2023; Pinyoanuntapong et al., 2024; Zhang et al., 2023b; Chen et al., 2023; Zhang et al., 2024). The existing methods of text-to-motion encompass several different approaches. For example, some utilize vector quantization and have achieved high generation performance (Zhang et al., 2023b; Guo et al., 2023; Lou et al., 2023; Pinyoanuntapong et al., 2024). Inspired by the development of diffusion models and their convenience, several diffusion-based text-to-motion methods have also been developed to not only generate motion but also to perform editing of the motion samples (Tevet et al., 2023; Zhang et al., 2024). In addition, Chen et al. adopted a latent diffusion model (Rombach et al., 2022) to enable fast motion generation (Chen et al., 2023). Alternatively, other methods for motion editing tasks utilize a pre-trained text-to-motion model as a generative prior, which has attached attention due to its applicability (Song et al., 2023b; Karunratanakul et al., 2023b; Xie et al., 2023; Karunratanakul et al., 2023a).
Ideally, (text-to-)motion generation models should be able to 1) perform fast generation, 2) achieve a high generation performance, and 3) perform editing motions. However, it is difficult for the existing approaches to meet all three requirements. The VQ-based text-to-motion methods (Zhang et al., 2023b; Guo et al., 2023; Lou et al., 2023; Pinyoanuntapong et al., 2024) require additional training for each editing task, as discussed in (Pinyoanuntapong et al., 2024). In other words, the VQ-based methods cannot control arbitrary joints in a training-free manner. Requiring extra training per editing task harms the applicability of motion generation models to production. Existing motion editing methods (Song et al., 2023b; Karunratanakul et al., 2023b; Xie et al., 2023; Karunratanakul et al., 2023a) utilize diffusion-based methods in the data-space, such as Motion Diffusion Model (MDM) (Tevet et al., 2023), as a generative prior, but the generation speed of diffusion in the data-space is not fast. In contrast, the existing latent diffusion-based motion generation method (Chen et al., 2023) performs fast generation, but motion editing based on continuous latent diffusion has not been explored yet.
In this paper, we aim to develop a text-to-motion model that achieves 1) fast generation, 2) a high generation performance, and 3) multiple motion editing in a training-free manner. We summarize the target properties satisfied by our model in Table 1 in comparison with the existing text-to-motion studies. To achieve this goal, we employ a motion variational autoencoder (VAE) to learn a low-dimensional latent representation for diverse human motion sequences and a latent diffusion model using text condition as in (Chen et al., 2023). Here, we combine a motion VAE with adversarial training to achieve better reconstruction/generation. It has been verified that adversarial training improves the performance of VAE in the image and audio domains (Esser et al., 2021; Iashin and Rahtu, 2021; Rombach et al., 2022; Liu et al., 2023), but not yet in motion generation. To the best of our knowledge, our model is the first latent-space motion generation model featuring a VAE is trained with adversarial training. In addition, we introduce training-free guided diffusion (He et al., 2024) into diffusion sampling on latent space to enable multiple motion editing tasks. We empirically show that adversarial training increases the quality on two widely utilized datasets. We also demonstrate that training-free guided diffusion can enable multiple motion editing tasks such as in-betweening, upper body editing, and path-following.
The main contributions of this paper are as follows:
@@
  • We propose a new text-to-motion model that can perform fast and high-quality generation and can deal with various types of motion editing tasks (motion in-betweening, upper body editing, and path following) in a training-free manner. To the best of our knowledge, this is the first study to demonstrate motion editing with latent diffusion.
  • We quantitatively show that introducing adversarial training to a motion VAE improves reconstruction and generation performance on text-to-motion. To the best of our knowledge, this is the first latent-space motion generation model that introduces adversarial training.
@@ 2. Related Work
2.1. Deep Generative Modeling
The progress in generative modeling has developed rapidly over the past decade (Kingma and Welling, 2013; Goodfellow et al., 2014; Sohl-Dickstein et al., 2015). In recent years, diffusion models have been actively explored, particularly in the image domain. They have achieved state-of-the-art performance not only in image generation (Gao et al., 2023; Peebles and Xie, 2023) but also in advanced generative tasks such as text-to-image (Saharia et al., 2022; Ramesh et al., 2022; Rombach et al., 2022), image inpainting (Lugmayr et al., 2022; Wang et al., 2022), and guided generation (Yu et al., 2023; Zhang et al., 2023a). A key approach for reducing inference costs is to reduce the essential dimensionality of the target data by learning its compressed representation (Rombach et al., 2022; Gao et al., 2023). This approach involves two stages: learning a VAE to acquire latent representations and then training a diffusion model on the low-dimensional representations.
A pioneering work that incorporates such a two-stage training scheme into generative modeling is VQ-VAE, which aims to compress the target data in a discrete manner (van den Oord et al., 2017). Several subsequent works have followed and expanded this discrete approach (Razavi et al., 2019; Esser et al., 2021; Lee et al., 2022). Esser et al. attempted to improve the reconstruction fidelity of VQ-VAE through adversarial training (Esser et al., 2021), which has been adopted by many subsequent works (Lee et al., 2022; Gu et al., 2022; Chang et al., 2022, 2023). As another advancement in the stage 1 model, Lee et al. generalized the VQ operation to residual vector quantization (RVQ), precisely approximating the encoded feature as a sequence of code stacks (Lee et al., 2022). In contrast, VQ-Diffusion adopts a discrete diffusion model that allows the input of a text prompt for learning the discrete latent, thereby achieving text-to-image generation (Gu et al., 2022). Mask-prediction training schemes with confidence-based iterative decoding have been proposed for the stage 2 (Chang et al., 2022, 2023).
2.2. Training-free Guided Generation
Starting with works such as (Song et al., 2021b; Meng et al., 2022), there have been many attempts to utilize pre-trained diffusion models as rich priors for various conditional generation tasks without additional training (i.e., training-free(Chung et al., 2023; Wang et al., 2022; Song et al., 2023b, a). The fundamental idea behind these approaches is to guide the intermediate samples of the diffusion model during the generation so that the final generated output satisfies the given conditions.
With increasing attention on the practical usefulness of latent diffusion models (LDMs), some recent works have focused on methods capable of dealing with LDMs. FreeDoM (Yu et al., 2023) performs a variety of conditional generation tasks in a unified framework by designing a task-specific loss function and utilizing the loss during guided sampling. Approaches such as PSLD (Rout et al., 2023) and MPGD (He et al., 2024) are attempting to improve on this even more, achieving higher performance and faster guided generation.
@@ 2.3. Motion Generation
Text-to-motion generation technology has made rapid progress with the development of neural networks (Ahuja and Morency, 2019; Petrovich et al., 2022; Guo et al., 2022; Tevet et al., 2022; Petrovich et al., 2023). In particular, diffusion models and VQ-based models have significantly advanced the field of motion generation.
MDM (Tevet et al., 2023) and MotionDiffuse (Zhang et al., 2024) adopt diffusion models for motion generation, which leads to better performance in terms of generation quality. However, these methods directly apply diffusion processes to raw motion sequences, thus resulting in slow generation. MLD (Chen et al., 2023) mitigates this issue by adopting a diffusion model in a low-dimensional latent space provided by a VAE trained on motion data, inspired by latent diffusion models (Rombach et al., 2022).
VQ-based models have been studied as well, inspired by the success of VQ-based models in image generation (Gu et al., 2022; Yu et al., 2022; Chang et al., 2023). In this approach, a VQ-VAE model is first trained on motion data to acquire discrete motion representations, and deep generative models are then applied to generate sequences of discrete representations (also called tokens). T2M-GPT (Zhang et al., 2023b), AttT2M (Zhong et al., 2023), and MotionGPT (Jiang et al., 2023) utilize autoregressive (AR) models to generate motion tokens. However, AR models are slow in inference because motion tokens are generated sequentially. To address this issue, M2DM (Kong et al., 2023) and DiverseMotion (Lou et al., 2023) apply discrete diffusion models to motion tokens in a latent space, whereas MMM (Pinyoanuntapong et al., 2024) adopts a mask prediction model. In another line of research, MoMask (Guo et al., 2023) leverages the RVQ technique to reduce errors caused by the VQ process.
In this paper, we adopt a continuous diffusion model to perform multiple editing tasks with a single framework (Section 3.4). To achieve high-fidelity and fast generation, we boost the performance of a VAE model with adversarial training and apply a diffusion model in the latent space provided by the trained VAE.
@@ 2.4. Motion Editing
Motion editing has attracted much research interest as well. MDM (Tevet et al., 2023) demonstrated upper body editing and motion in-betweening by applying diffusion inpainting to motion data in both the spatial and temporal domains. LGD (Song et al., 2023b) demonstrated path-following motion generation with a guided diffusion that utilizes multiple samples from a suitable distribution to reduce bias. GMD (Karunratanakul et al., 2023b) guides the position of the root joint to control motion trajectories. OmniControl (Xie et al., 2023) controls any joints at any time by guiding a pre-trained motion diffusion model with an analytic function. DNO (Karunratanakul et al., 2023a) optimizes the diffusion latent noise of a pre-trained text-to-motion model with user-provided criteria in the motion space and achieves multiple editing tasks. However, these methods employ data-space diffusion models similar to MDM and have not been demonstrated with a latent diffusion model. Recently, MMM (Pinyoanuntapong et al., 2024) demonstrated motion editing by placing masked tokens in the place that needs editing and applying the mask prediction framework. However, it cannot control an arbitrary body part without additional model training, and the authors trained separate upper-body and lower-body generative models to perform upper-body editing.
In this paper, we employ a latent diffusion model to achieve fast generation and apply a training-free guided generation framework that can deal with multiple editing tasks without additional model training.
@@ 3. Method
3.1. Overall Framework of MoLA
The goal of this study is to develop a framework for fast and high-quality text-guided motion generation and to deal with multiple control tasks in a training-free way. To achieve this, we propose the following training and inference methods:
  • Stage 1 training: We first train a motion VAE enhanced by adversarial training to achieve high reconstruction/generation performance. We further propose replacing the decoder for adversarial training with a SAN-based one to boost our model performance (Section 3.2).
  • Stage 2 training: To reduce computational complexity while simultaneously enabling high-quality text-driven motion generation, we train a text-conditioned diffusion model on the low-dimensional latent space learned by stage 1, i.e., a latent diffusion model (Section 3.3).
  • Guided generation: Adopting training-free guided generation in inference enables multiple editing functions required in motion generation without additional training (Section 3.4).
The outline of our text-to-motion generation model, MoLA, is shown in Figure 2.
Figure 2. Overall architecture of MoLA. Stage 1: A motion VAE enhanced by adversarial training learns a low-dimensional latent representation for diverse human motion sequences. Stage 2: A text-conditioned latent diffusion model reduces computational costs and enables high-quality text-to-motion generation by utilizing the latent representation obtained in stage 1. Guided generation: During the inference of the stage 2 model, a training-free guided generation method minimizes a loss function $L$, which can be flexibly set depending on the task, and multiple motion editing tasks can also be performed in a single framework.
@@ 3.2. Stage 1: Continuous Motion Latent Representation with Adversarial Training
3.2.1. Learning a rich motion latent representation with VAE-GAN
We propose a motion variational autoencoder (VAE), enhanced by adversarial training, to learn a low-dimensional latent representation for diverse human motion sequences. Assume an observed motion $\bm{x}\in\mathbb{R}^{N\times L}$, where $N$ and $L$ denote raw motion data dimension per frame and motion length, respectively. To learn such a latent representation, we first define a $d_{z}$-dimensional latent variable $\bm{z}\in\mathbb{R}^{d_{z}}$, which is assumed to generate data sample $\bm{x}$. The generative process is modeled as $\bm{x}\sim p_{\bm{\psi}}(\bm{x}|\bm{z})$ with a prior $p(\bm{z})$. The prior is assumed to be a standard Gaussian distribution, i.e., $p(\bm{z})=\mathcal{N}(\bm{0},\bm{I})$. We model the conditional distribution as a Gaussian distribution: $p_{\bm{\psi}}(\bm{x}|\bm{z})=\mathcal{N}(g_{\bm{\psi}}(\bm{z}),\sigma^{2}\bm{I})$ with $g_{\bm{\psi}}:\mathbb{R}^{d_{z}}\to\mathbb{R}^{N\times L}$ and $\sigma^{2}\in\mathbb{R}_{+}$, where $\mathbb{R}_{+}$ indicates the set of all positive real numbers. As in a usual VAE, we introduce an approximated posterior, which is modeled by a Gaussian distribution as $q_{\bm{\eta}}(\bm{z}|\bm{x})=\mathcal{N}(\bm{\mu}_{\bm{\eta}}(\bm{x}),\mathrm{diag}(\bm{\sigma}_{\bm{\eta}}(\bm{x})))$, with $\bm{\mu}_{\bm{\eta}}:\mathbb{R}^{N\times L}\to\mathbb{R}^{d_{z}}$ and $\bm{\sigma}_{\bm{\eta}}:\mathbb{R}^{N\times L}\to\mathbb{R}_{+}^{d_{z}}$. As a result, the VAE consists of an encoder and a decoder, parameterized by $\bm{\eta}$ and $\bm{\psi}$, respectively. We construct the autoencoder based on a transformer-based architecture (Petrovich et al., 2021; Chen et al., 2023). The objective function for the VAE is formulated as the negative evidence lower bound (negative ELBO) per sample $\bm{x}$, which is a summation of mean squared error and KL regularization terms:

### v3

@@ Abstract.
In text-to-motion generation, controllability as well as generation quality and speed has become increasingly critical. The controllability challenges include generating a motion of a length that matches the given textual description and editing the generated motions according to control signals, such as the start-end positions and the pelvis trajectory. In this paper, we propose MoLA, which provides fast, high-quality, variable-length motion generation and can also deal with multiple editing tasks in a single framework. Our approach revisits the motion representation used as inputs and outputs in the model, incorporating an activation variable to enable variable-length motion generation. Additionally, we integrate a variational autoencoder and a latent diffusion model, further enhanced through adversarial training, to achieve high-quality and fast generation. Moreover, we apply a training-free guided generation framework to achieve various editing tasks with motion control inputs. We quantitatively show the effectiveness of adversarial learning in text-to-motion generation, and demonstrate the applicability of our editing framework to multiple editing tasks in the motion domain. Our code is available at https://github.com/sony/MoLA
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: XXXjournalvolume: XXXjournalnumber: XXXarticle: XXXccs: Computing methodologies Computer visionccs: Computing methodologies Motion processing
Figure 1. MoLA achieves fast and high-quality human motion generation given textual descriptions while enabling motion editing applications. With MoLA, we can deal with various types of motion editing tasks in a single framework.
@@ 1. Introduction
Human motion synthesis from text is an emerging task with highly relevant applications in fields such as multimedia production and computer animation. For example, an animator might wish to create or edit a motion prototype to verify their artistic intent before time-consuming animation or motion capture commences, generate smooth in-between motion between two motion capture clips, or generate specific motion that follows a predefined trajectory. In Figure 1, we demonstrate results of our approach on such motion generation and editing tasks. To be useful in those real-world applications, a method has to excel in three domains: (1) motion quality, which encompasses both the general motion quality and the adherence to the textual description; (2) fast inference time; (3) easy motion editing.
Several recent works have attempted to address these desired properties: Methods based on vector quantization (VQ), such as T2M-GPT (Zhang et al., 2023b), MoMask (Guo et al., 2024), MMM (Pinyoanuntapong et al., 2024b), ParCo (Zou et al., 2024), and BAMM (Pinyoanuntapong et al., 2024a), achieve impressive generation quality by compressing human motion into discrete tokens and then sampling those tokens to synthesize motion. Diffusion-based methods in data space, such as MDM (Tevet et al., 2023), provide impressive flexibility with regard to quality and motion editability. To further improve motion quality and inference time, latent-space based methods such as MLD (Chen et al., 2023) or MotionMamba (Zhang et al., 2025) operate in a learned continuous latent space.
However, no state-of-the-art method excels in all three domains necessary for real-world applications. For example, while latent-space based approaches such as VQ-based methods or MLD achieve impressive motion quality and fast inference time, they cannot edit a given motion sequence in a training-free manner. In contrast, data-space based methods such as MDM are able to edit motion in a training-free manner, which, however, comes at the cost of slow inference time and lower generation quality. Moreover, while these models excel at motion generation, they require manual control of the generated motion length, which limits their flexibility. For instance, MoMask (Guo et al., 2024) uses a length estimator during inference to predict motion length based on text input. However, inaccurate predictions may result in significant motion drift (Wu et al., 2024).
To close this gap, we propose MoLA, Motion Generation and Editing with Latent Diffusion Enhanced by Adversarial Training. We revisit the representation of motion features used as inputs and outputs of the model and introduce an activation variable that characterizes the length of the motion. We utilize a variational auto-encoder (VAE) (Kingma and Welling, 2013) for its continuous latent space, which allows training-free editing. This is in contrast to discrete latent spaces that do not allow for training-free motion editing. To bridge the motion quality gap, we enhance the VAE training with adversarial learning, which has been shown to work well in the image (Esser et al., 2021; Iashin and Rahtu, 2021; Rombach et al., 2022) and audio (Liu et al., 2023) domains. We empirically show that our model achieves variable-length motion generation aligned with textual descriptions and high generation performance on a commonly used dataset. We also demonstrate that training-free guided diffusion can enable multiple motion editing tasks such as path-following, in-betweening, and upper body editing. These experiments also highlight the significant improvement in speed and performance of our model compared to the performance of existing training-free motion editing models (see Figure 2).
@@
The main contributions of this paper are threefold. First, we introduce an activation variable into the motion representation and show that our method can perform variable-length motion generation conditioned on text. Second, we propose a new continuous latent-based motion generation model that introduces adversarial training into the motion VAE and quantitatively show that it significantly pushes the limits of existing continuous-based methods. Finally, we demonstrate that our model not only shows high generation performance but also can deal with various types of motion editing tasks (path following, motion in-betweening, and upper body editing) in a training-free manner.
Figure 2. Comparison of inference cost, generation performance, and editability for text-to-motion methods. $\bullet$ means a method that can edit motion in a training-free manner, and $\times$ means a method that cannot edit motion in a training-free manner. We calculate the average inference time on the test set of HumanML3D without model or data loading parts. All tests are performed on the same NVIDIA A100 GPU. The pink arrow in the figure indicates that our method significantly extends the performance boundaries (in terms of generation quality and speed) of methods categorized as enabling training-free editing.
@@ 2.1. Motion Generation
Text-to-motion generation technology has made rapid progress with the diffusion models and VQ-based models. MDM (Tevet et al., 2023) and MotionDiffuse (Zhang et al., 2024) adopt diffusion models for motion generation, which leads to better performance in terms of generation quality. However, these methods directly apply diffusion processes to raw motion sequences, thus resulting in slow generation. MLD (Chen et al., 2023) mitigates this issue by adopting a diffusion model in a low-dimensional latent space provided by a VAE trained on motion data, inspired by latent diffusion models (Rombach et al., 2022). VQ-based models have been studied as well, inspired by the success of VQ-based models in image generation (Gu et al., 2022; Yu et al., 2022; Chang et al., 2023). In this approach, a VQ-VAE model is first trained on motion data to acquire discrete motion representations, and deep generative models are then applied to generate sequences of discrete representations (also called tokens). T2M-GPT (Zhang et al., 2023b), AttT2M (Zhong et al., 2023), MotionGPT (Jiang et al., 2023) and ParCo (Zou et al., 2024) utilize autoregressive (AR) models to generate motion tokens. However, AR models are slow in inference because motion tokens are generated sequentially. To address this issue, M2DM (Kong et al., 2023) and DiverseMotion (Lou et al., 2023) apply discrete diffusion models to motion tokens in a latent space, whereas MMM (Pinyoanuntapong et al., 2024b), MoMask (Guo et al., 2024) and BAMM (Pinyoanuntapong et al., 2024a) adopt a mask prediction model.
Figure 3. The overall framework of MoLA. Stage 1: A motion VAE enhanced by adversarial training learns a low-dimensional latent representation of diverse motion sequences. Stage 2: A text-conditioned latent diffusion model leverages this representation for fast and high-quality text-to-motion generation. Guided generation: During inference, a gradient-based method minimizes a loss function $\mathfrak{L}_{\text{Motion}}$ for each desired editing task, enabling multiple motion editing tasks within a unified framework.
@@ 2.2. Motion Editing
Motion editing has attracted much research interest as well. MDM (Tevet et al., 2023) demonstrated upper body editing and motion in-betweening by applying diffusion inpainting to motion data in both the spatial and temporal domains. LGD (Song et al., 2023) demonstrated path-following motion generation with a guided diffusion that utilizes multiple samples from a suitable distribution to reduce bias. GMD (Karunratanakul et al., 2023) guides the position of the root joint to control motion trajectories. OmniControl (Xie et al., 2024) controls any joints at any time by guiding a pre-trained motion diffusion model with an analytic function. DNO (Karunratanakul et al., 2024) optimizes the diffusion latent noise of a pre-trained text-to-motion model with user-provided criteria in the motion space and achieves multiple editing tasks. However, these methods employ data-space diffusion models similar to MDM and have not been demonstrated with a latent diffusion model. Recently, MMM (Pinyoanuntapong et al., 2024b) demonstrated motion editing by placing masked tokens in the place that needs editing and applying the mask prediction framework, and MotionLCM (Dai et al., 2024) proposed a fast controllable motion generation framework by introducing latent consistency distillation and the motion ControlNet (Zhang et al., 2023a) manipulation in the latent space. However, their methods cannot control an arbitrary body part in a training-free manner.
@@ 3. Method
The goal of this study is to develop a framework for fast and high-quality text-guided motion generation and to deal with multiple control tasks in a training-free way. To achieve this, we propose the following training and inference tricks (I)-(IV): (I) We review the representation of motion features to achieve improved accuracy and variable length motion generation. We select the necessary features to reduce the burden on our model’s VAE encoder and add an activation variable to the representation to determine the motion length (Section 3.1). (II) We train a motion VAE enhanced by adversarial training to achieve high generation performance (Section 3.2). (III) To reduce computational complexity while simultaneously enabling high-quality text-driven motion generation, we train a text-conditioned diffusion model on the low-dimensional latent space obtained by VAE model (Section 3.3). (IV) Guided generation: Adopting training-free guided generation in inference enables multiple editing functions required in motion generation without additional training (Section 3.4). The outline of our text-to-motion generation model, MoLA, is shown in Figure 3.
3.1. Motion representation
We build upon the motion representation used in (Guo et al., 2022; Plappert et al., 2016) and improve the representation to enable variable length generation and improve performance in our framework. The pose representation used in (Guo et al., 2022; Plappert et al., 2016) is expressed as $\bm{m}\in\mathbb{R}^{(4+12N_{j}+4)\times L}$, where $N_{j}$ is the number of joints. A motion can be represented as a sequence of this representation. The $i$-th $(1\leq i\leq L)$ pose is defined as follows:
(1) $\bm{m}^{i}=[\dot{r}_{a}^{i},\dot{r}_{x}^{i},\dot{r}_{z}^{i},r_{y}^{i},\bm{j}_{p}^{i},\bm{j}_{r}^{i},\bm{j}_{v}^{i},\bm{c}^{i}]$
where $\dot{r}_{a}\in\mathbb{R}$ is root angular velocity along the $Y$-axis, $\dot{r}_{x}$ and $\dot{r}_{z}\in\mathbb{R}$ are root linear velocities on $XZ$-plane, $r_{y}$ is root height, $\bm{j}_{p}\in\mathbb{R}^{3(N_{j}-1)}$ is local joints positions, $\bm{j}_{r}\in\mathbb{R}^{6(N_{j}-1)}$ is rotations in root space, $\bm{j}_{v}\in\mathbb{R}^{3N_{j}}$ is velocities and $\bm{c}\in\mathbb{R}^{4}$ is binary foot-ground contact features by thresholding the heel and toe joint velocities. To achieve variable-length motion generation during stage 2 inference (explained in Section 3.3), we concatenate an activation variable $a^{i}\in\mathbb{R}$ for the $i$-th pose to Equation (1) as $[\bm{m}^{i},a^{i}]$. Additionally, to reduce the burden on the encoder in Section 3.2, we remove the substantially redundant information $\bm{j}_{v}$ and $\bm{c}$ from Equation (1) as $\tilde{\bm{m}}^{i}=[\dot{r}_{a}^{i},\dot{r}_{x}^{i},\dot{r}_{z}^{i},r_{y}^{i},\bm{j}_{p}^{i},\bm{c}^{i}]$, and then we set the vector for the encoder input in Section 3.2 as follows:
(2) $\bm{x}^{i}=[\tilde{\bm{m}}^{i},a^{i}]=[\dot{r}_{a}^{i},\dot{r}_{x}^{i},\dot{r}_{z}^{i},r_{y}^{i},\bm{j}_{p}^{i},\bm{j}_{r}^{i},a^{i}].$
[…]
@@ 3.2. Stage 1: Continuous Motion Latent Representation with Adversarial Training
3.2.1. Learning a motion latent representation with VAE-GAN
We propose a motion variational autoencoder (VAE), enhanced by adversarial training, to learn a low-dimensional latent representation for diverse human motion sequences. Assume an observed motion $\bm{x}\in\mathbb{R}^{N\times L}$, where $N$ and $L$ denote raw motion data dimension per frame and motion length, respectively. To learn such a latent representation, we first define a latent variable $\bm{z}\in\mathbb{R}^{d_{z}\times d_{l}}$, which is assumed to generate data sample $\bm{x}$, where $d_{z},d_{l}\in\mathbb{N}$. The generative process is modeled as $\bm{x}\sim p_{\bm{\psi}}(\bm{x}|\bm{z})$ with a prior $p(\bm{z})$. The prior is assumed to be a standard Gaussian distribution, i.e., $p(\bm{z})=\mathcal{N}(\bm{0},\bm{I})$. We model the conditional distribution as a Gaussian distribution: $p_{\bm{\psi}}(\bm{x}|\bm{z})=\mathcal{N}(g_{\bm{\psi}}(\bm{z}),\sigma^{2}\bm{I})$ with $g_{\bm{\psi}}:\mathbb{R}^{d_{z}\times d_{l}}\to\mathbb{R}^{N\times L}$ and $\sigma^{2}\in\mathbb{R}_{+}$, where $\mathbb{R}_{+}$ indicates the set of all positive real numbers. As in a usual VAE, we introduce an approximated posterior, which is modeled by $q_{\bm{\eta}}(\bm{z}|\bm{x}):\mathbb{R}^{N\times L}\to\mathbb{R}^{d_{z}\times d_{l}}$. As a result, the VAE consists of an encoder and a decoder, parameterized by $\bm{\eta}$ and $\bm{\psi}$, respectively. The objective function for the VAE is formulated as the negative evidence lower bound (negative ELBO) per sample $\bm{x}$, which is a weighted summation of mean squared error and KL regularization terms: